OpenStudy (cimbie):

What wavelength of light would be required to excite the electron in a ground state hydrogen atom to an orbital where n=5?

OpenStudy (jiteshmeghwal9):

(1/x)=109677((1/n_1)-(1/n_2))

OpenStudy (cuanchi):

95nm

OpenStudy (jiteshmeghwal9):

X=wavelength N_1=initial orbital N_2=final orbital

OpenStudy (jiteshmeghwal9):

(1/x)=109677(1/1-1/5)

OpenStudy (jiteshmeghwal9):

Solve for x

OpenStudy (cuanchi):

RH is 1.09678 x 10-2 nm-1 n1 and n2 has to be (n1)^2 and (n2)^2